OpenStudy (anonymous):

Tiana can clear a football field of debris in 3 hours. Jacob can clear the same field in 2 hours. If they decide to clear the field together, how long will it take them? ~ 50 minutes ~ 1 hour 5 minutes ~ 1 hour 12 minutes ~ 5 hours

OpenStudy (anonymous):

Tianas debris clearing rate is 1/3 football fields per hour and Jacob´s 1/2. 1/3*t+1/2*t=1

OpenStudy (anonymous):

1/3*t+1/2*t=1 Combine like terms 5/6*t=1 Divide both sides by 5/6 t=1.2 1.2 hours = 1 hour and 12 mins. Correct?